Find the critical value, tc, for c = 0.95 and n= 16. O A. 2.947 OB. 2.131 OC. 2.120 OD. 2.602
Answer #1

Solution :

n = 16

Degrees of freedom = df = n - 1 = 16 - 1 = 15

c = 0.95

At 95% confidence level the t is ,

\alpha = 1 - c = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha/2,df = t0.025,15 = 2.131 using excel ( = TINV(probability, degreed of freedom)

Critical value = 2.131

Ans: option B) 2.131 is correct